OpenStudy (anonymous):

A 4400 W motor is used to do work. If the motor is used for 200 s, how much work could it do? (Power: P = W/t) 22 J 4200 J 4600 J 880,000 J

OpenStudy (johnweldon1993):

Well we have the equation given \[\large Power = \frac{Work}{Time}\] So that means \[\large Work = Power \times Time\] We see that Power = 4400W And we see that Times = 200s So \[\large Work = 4400W \times 200s = ?\]

OpenStudy (anonymous):

880000?

OpenStudy (johnweldon1993):

Correct \[\large Work = 880,000Ws = 880,000J\]

OpenStudy (anonymous):

Thank you!!

OpenStudy (johnweldon1993):

No problem!

OpenStudy (anonymous):

How about this one? In the reaction C + O2 → CO2, 18 g of carbon react with oxygen to produce 72 g of carbon dioxide. What mass of oxygen would be needed in the reaction? 18 g 54 g 72 g 90 g

OpenStudy (johnweldon1993):

Sorry not good at all with chemistry! :(

OpenStudy (anonymous):

It's okay neither am I! ;D